NPP Projection Report 2/19 Close


To do this evening/ was completed: In addition to Marty and Moment, we also have no more shows of Secret Agent and No Other Choice starting tomorrow, so these posters/marquee letters should get taken down.

Presentation Issues: 

Preshow Clips/Playlist: No lights down cue in any version of the Sirat playlist. I added for the 2/20 ones since the 2/13 playlist was only being used for the one show tonight. Manually brought them down for tonight's show.

Trailer Updates:

Special Events/ AV set up: American Cyborg: Steel Warrior went well. I didn't see the motor start cue on the second reel so the audience saw a bit of leader. I checked the print at the bench after taking it off the projector and the motor cue was there, but out of frame for 1.85 so it did not appear on screen. Try and check for this when inspecting, for 1.85 the cue should be right by the second sprocket from the top of the frame.

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): DCP'd the last two Dark Room shorts and ingested them to the LMS.

Scream 7 (😞) ingested to the LMS

Outgoing/Shipping (content): John got HFC to send us the return labels for Ghost, Snow Day, and Pulp Fiction. These are all Paramount prints. Printed all three and scheduled a pickup for Ghost, it's on the cart. Labels for the other two are in their respective cans.

Printed labels and scheduled a pickup for American Cyborg, this is also on the cart.

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning:

Cue sheets updated/checked (every shift): Yes, all set for tomorrow.

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will

NPP Projection Admin 7/6